OpenStudy (anonymous):

x^2 – 2x = 3? a. x = 4 and x = –1 b. x = –3 and x = 1 c. x = 4 and x = –3 d. x = 3 and x = –1

OpenStudy (anonymous):

For equations level 2 (exponential 2), you can solve the equations for the follow equation: \[\frac{ -b \pm \sqrt{b ^{2}- 4ac} }{ 2a }= x \] when a is the coeficient of x^2, b is the coeficient of x and c is the independient variable (in your case a=1, b=-2,c=-3)

OpenStudy (anonymous):

this formule give to you 2 answer when \[\frac{ -b + \sqrt{b ^{2}- 4ac} }{ 2a }\] and when \[\frac{ -b - \sqrt{b ^{2}- 4ac} }{ 2a }\]

OpenStudy (anonymous):

is it 4 and -1?

OpenStudy (anonymous):

i think that is d.. only replace a, b and c.. you get the answer so ;)

OpenStudy (amoodarya):

x^2-2x-3=0 look at this file